# Exploring the spectral characteristics of the periodic burster 4U 1323-62: Type-I X-ray burst and persistent emission

### v2

@@ Abstract
We report on the results obtained by the analysis of persistent and type-I thermonuclear X-ray burst emission observed from the periodic burster 4U 1323-62. These analyses are based on the NuSTAR observation performed on 2024 August 7 for a total exposure of around 90 ks. The persistent emission is well described by an absorbed thermal Comptonization model. An absorption edge is also detected at an energy of 7.87 keV, which indicates the presence of absorbing material in the vicinity of this system. Six bursts have been observed during this observation, wherein we find the burst recurrence time to be hr. All the bursts exhibit the characteristics of a sharp rise and exponential decay. We perform the time-resolved spectroscopy of the burst spectra described by a model consisting of thermal emission from the neutron star surface and a varying persistent emission component to study the evolution of burst parameters. The enhancement of the persistent emission during burst exposure is characterized by the scaling parameter , which reflects the increasing strength of the burst–disc interaction with burst intensity, likely driven by Poynting–Robertson drag. The spectral analysis of bursts estimate the average apparent blackbody emitting radius of the neutron star to lie within 1.5–3.5 km. The ignition depths computed from the burst parameters indicate short Type-I thermonuclear bursts from a mixed hydrogen-helium fuel layer.
@@ 3 Light Curve
The rise times of bursts B1–B6 were measured to be 7 s, 4 s, 7 s, 2 s, 7 s, and 4 s, respectively, while their decay tails were significantly longer—95 s, 102 s, 101 s, 33 s, 99 s, and 83 s. To examine their spectral evolution, we conducted time-resolved spectral analysis for all bursts. Each of the bursts, barring B4 and B6, have exposure above 100 s. For B4 and B6, the burst exposures were shorter (36s and 91s, respectively). A time segment of 20 ks duration, from elapsed time 30th ks to 50th ks, was used to ascertain the flux parameter of persistent emission.
@@ 4.1 Persistent spectral analysis
We extracted the persistent spectrum of the source from a 20 ks time interval (from the elapsed time of 30th ks to 50th ks) of the current observation using the NuSTAR FPMA and FPMB data. No X-ray bursts and prominent intensity dips have been observed in this time interval. We fit them simultaneously in the energy band 3 to 60 keV, where the source counts are well above the background counts. We include a constant multiplication factor to account for cross-calibration of the two instruments, FPMA and FPMB.
@@
Initially, we attempted to fit the persistent spectrum with an absorbed cutoff power law model (Tbabs*cutoffpl), which is referred to as Model 1. The Model 1 resulted in = 599/518 with a photon index of = 1.50 0.02 and the cutoff energy Ecut = 80 keV. Previous studies of this source also reported that the energy spectrum is dominated by a power-law component. The measured parameter values, and Ecut, are consistent with values obtained from BeppoSAX and Suzaku observations (boirin2005highly; balucinska2009neutral).
@@
To describe the persistent emission more precisely, we employed a thermal Comptonization model, nthcomp (1996MNRAS.283..193Z; 1999MNRAS.309..561Z), which involves a seed photon source as a blackbody or a disk-like multi-color blackbody. These seed photons are up-scattered by hot electrons, resulting in the output spectrum. We assumed that the source of the seed photons is the accretion disk ( inptype = 1). This model Tbabs*nthcomp (Model 2) resulted in = 550/517 ( = -49 for 1 additional parameter, as compared to Model 1) with a photon index of , the electron temperature kTe=31 keV, and the seed temperature kTbb = 1.31 0.06 keV. The electron temperature of the corona is comparable to the observed cut-off energy of the spectrum. In addition, we found an absorption feature at 7-8 keV and tried to fit it with an absorption edge model. The inclusion of the edge model, Tbabs*edge*nthcomp (Model 3), improved the fit significantly to = 511/515 ( = -39 for 2 additional parameters, compared to Model 2) with edge energy E = 7.87 0.11 keV and optical depth = 0.12 0.02.
@@
To provide a more accurate model description, we further replaced nthcomp with the convolution model thcomp (1996MNRAS.283..193Z; 2019MNRAS.485.2942N; 2020MNRAS.492.5234Z), an updated version of nthcomp that aligns more closely with the actual Monte Carlo spectra from Comptonization than nthcomp does. The model thcomp involves the upscattering fraction (fsc) of photons from any seed emission and gives the low-energy cut-off and the high-energy rollover corresponding to the soft tail of the seed photon and electron temperature of the corona, respectively. As we assumed that the accretion disk supplies the seed photons, we applied the model Tbabs*edge*thcomp*diskbb (Model 4). This model also satisfactorily described the persistent spectrum, with = 506/514, and provided similar best-fit parameter values, including absorption edge energies compared to Model 3, suggesting that the source geometry is correctly identified. All the best-fit parameters for the persistent spectral fitting are listed in Table 2. The persistent spectra for the best fit values corresponding to Model 3 is shown in Fig. 3.
@@
We also explored the pre-burst persistent spectra of the source before examining the X-ray burst emissions in detail. For this, we extracted the spectra of the persistent emission from exposures of 600, 600, 1000, 300, 1000, and 1000 s, respectively, just before the onset of bursts B1, B2, B3, B4, B5, and B6, depending on the data availability on the same NuSTAR orbit. The 3 – 60 keV energy spectrum of the pre-burst emission was also successfully fitted with an absorbed thermal Comptonization model along with an absorption edge (Model 3). We note that the best-fit parameter values are similar within error when comparing the values obtained from the persistent emission of the 20-ks exposures described earlier. However, parameters such as electron temperature (k) and edge energy (Eedge) are not well constrained by the pre-burst persistent spectra due to the limited exposures and low statistics (low count rate) before X-ray bursts. We fixed them to the best-fit values obtained earlier. The power-law photon index , seed photon temperature k, absorption optical depth , and the corresponding values obtained from modeling the pre-burst emissions are reported in Table 3.
@@ 4.2 Burst spectral analysis
The burst spectra are shown in Fig. 4. As a best-fit value for all the bursts from Model 6, we obtained the blackbody temperature as k = 1.802 0.027 keV and radius for the blackbody emitting region km. The nthcomp norm is obtained as . The time-resolved spectral analysis was performed following a similar approach to that adopted by mondal2025nustar for the “clocked burster” GS 1826–24.
@@ 4.3 Time-resolved analysis of the burst spectra
The best-fit parameters of the B1 are shown in Table 4, and the corresponding spectra are shown in lower panel of Fig. 5. The blackbody emitting radius of the neutron star reaches a maximum 3.67 km during S2, where the bbodyrad flux peaks at ergs cm -2 s -1 and then gradually decreases. However, for some segments we obtained a low reduced value because of insufficient count rates.The blackbody temperature (k) for B1 reaches a maximum at S2 2.04 keV and then decreases monotonically till S5. The best-fit parameters for time-resolved spectroscopy of B2 are listed in Table 5 and the corresponding spectra are shown in the lower panel of Fig. 6. The bbodyrad flux reaches its peak value during segment S2, after which the bbodyrad flux steadily decreases through S5. The apparent blackbody emitting radius of the neutron star is estimated around 3.09 km during S2, increasing slightly to 3.29 km in S3. A larger radius of 3.71 km is obtained for S5; however, this value is poorly constrained due to limited statistics (0.696), and thus is not considered reliable. The evolution of the blackbody temperature () shows a maximum of keV in S2, followed by a gradual decrease toward S5. The best-fit spectral parameters for B5 across its five time segments are presented in Table 6, with the corresponding spectra and residuals shown in the lower panel of Fig. 7. The blackbody flux reaches its maximum during segment S2 ( erg cm -2 s -1) and subsequently declines in the later segments. The evolution of the blackbody radius closely follows the flux variation, attaining its largest value of approximately 3.5 km during S2, while the blackbody temperature () also attains its highest value at this segment. Across all bursts, the persistent emission shows a pronounced enhancement (), ranging within 1–4 in B1, 1–8 in B2, and 1–7 in B3, suggesting a variable mass accretion rate during the burst evolution (worpel2013evidence).
@@ 5 Discussion
We report on the NuSTAR observation of the periodic burster 4U 1323-62, performed on 2024 August 7, for a total exposure of 90 ks. We observed the presence of six thermonuclear X-ray bursts and regular intensity dips in the NuSTAR light curve of this source. The observed burst profiles are very similar to those extracted earlier from the XMM-Newton EPIC PN and AstroSat LAXPC detectors during a persistent interval (boirin2005highly; Bhulla_2020). The observed period ( 2.91-2.98 hr) of the intensity dips is also consistent with the other observations (1985SSRv...40..287V; boirin2005highly; Bhulla_2020). During this observation, the persistent emission is well described by the absorbed thermal Comptonization model in the energy band of 3–60 keV. Our persistent spectral analysis suggests that the Comptonization of soft photons originating from the accretion disc is occurring. The power law index , electron temperature of the corona k, and the seed photon temperatures kTbb obtained from the Comptonization model are 1.72 0.01, 23 4 keV, and keV, respectively. In addition, we detected an absorption edge at an energy of 7.87 keV, indicating the presence of an absorbing medium in the vicinity of the system. Apart from the dip intervals, the persistent flux remains unchanged on our observed timescale. We measured a 3-60 keV persistent flux of F erg cm-2 s-1, which corresponds to a persistent luminosity of Lper 3.321036 erg s-1 for an assumed distance of 10 kpc.
@@
distinct properties with , the shortest decay time (29.5 s), and the lowest fluence (0.385 erg cm-2). Such a high typically indicates either accretion-driven emission or incomplete fuel burning (10.1093/mnras/staa2749). In contrast, the subsequent bursts (B5 and B6) with 64–74 are consistent with mixed H/He-fueled events.
@@ 7 Acknowledgements
We thank NuSTARDAS for ensuring proper extraction of usable data. For the analysis of archival data we deeply acknowledge HEASOFT and CALDB. A token of thanks is extended towards the non-NET fellowship program of Visva-Bharati University for funding the research. We thank IUCAA for its vast repository of knowledge that enlightens curious minds.

### v3

@@ Abstract
We report on the results obtained by the analysis of persistent and type-I thermonuclear X-ray burst emission observed from the periodic burster 4U 1323-62. These analyses are based on the NuSTAR observation performed on 2024 August 7 for a total exposure of around 90 ks. The persistent emission is well described by an absorbed thermal Comptonization model. An absorption edge is also detected at an energy of 7.42 keV, which indicates the presence of absorbing material in the vicinity of this system. Six bursts have been observed during this observation, wherein we find the burst recurrence time to be hr. All the bursts exhibit the characteristics of a sharp rise and exponential decay. We perform the time-resolved spectroscopy of the burst spectra described by a model consisting of thermal emission from the neutron star surface and a varying persistent emission component to study the evolution of burst parameters. The enhancement of the persistent emission during burst exposure is characterized by the scaling parameter , which reflects the increasing strength of the burst–disc interaction with burst intensity, likely driven by Poynting–Robertson drag. The spectral analysis of bursts estimate the average apparent blackbody emitting radius of the neutron star to lie within 1.5–3.5 km. The ignition depths computed from the burst parameters indicate short Type-I thermonuclear bursts from a mixed hydrogen-helium fuel layer.
@@ 3 Light Curve
The rise times of bursts B1–B6 were measured to be 7 s, 4 s, 7 s, 2 s, 7 s, and 4 s, respectively, while their decay tails were significantly longer—95 s, 102 s, 101 s, 33 s, 99 s, and 83 s. To examine their spectral evolution, we conducted time-resolved spectral analysis for all bursts. Each of the bursts, barring B4 and B6, have exposure above 100 s. For B4 and B6, the burst exposures were shorter (36s and 91s, respectively). We also extracted the light curves for all bursts with a bin size of 0.5 s for the 30-70 keV energy band, starting from the onset of each burst. The light curves in the 3–70 keV (10 s bin size) and 30–70 keV energy ranges are plotted together for comparison and shown in the lower panel of Figure 2. From the inset plot in the lower panel of Fig. 2, the burst count rate in 30-70 keV energy band is approximately 4 counts/s at the peak and decreases to about 3 counts/s during the decay phase of the burst. An exposure of 45 ks, excluding the duration of bursts and dips, was used to extract the persistent spectra.
@@ 4.1 Persistent spectral analysis
We extracted the persistent spectrum of the source for a 45 ks duration of the current observation using the NuSTAR FPMA and FPMB. No X-ray bursts and prominent intensity dips have been observed in this time interval. We fit them simultaneously in the energy band 3 to 60 keV, where the source counts are well above the background counts. We include a constant multiplication factor to account for cross-calibration of the two instruments, FPMA and FPMB.
@@
Initially, we attempted to fit the persistent spectrum with an absorbed cutoff power law model (const*Tbabs*cutoffpl), which is referred to as Model 1. The Model 1 resulted in = 1222/881 with a photon index of = 1.46 0.01 and the cutoff energy Ecut = 72 keV. Previous studies of this source also reported that the energy spectrum is dominated by a power-law component. The measured parameter values, and Ecut, are consistent with values obtained from BeppoSAX and Suzaku observations (boirin2005highly; balucinska2009neutral).
@@
To describe the persistent emission more precisely, we employed a thermal Comptonization model, nthcomp (1996MNRAS.283..193Z; 1999MNRAS.309..561Z), which involves a seed photon source as a blackbody or a disk-like multi-color blackbody. These seed photons are up-scattered by hot electrons, resulting in the output spectrum. We assumed that the source of the seed photons is the accretion disk ( inptype = 1). This model const*Tbabs*nthcomp (Model 2) resulted in = 1047/880 ( = -175 for 1 additional parameter, as compared to Model 1) with a photon index of , the electron temperature kTe = keV, and the seed temperature kTbb = 1.33 0.03 keV. The electron temperature of the corona is comparable to the observed cut-off energy of the spectrum. In addition, we found an absorption feature at 7-8 keV and tried to fit it with an absorption edge model. The inclusion of the edge model, const*Tbabs*edge*nthcomp (Model 3), improved the fit significantly to = 928/878 ( = -119 for 2 additional parameters, compared to Model 2) with edge energy E = 7.42 keV and optical depth = 0.12 0.02. The absorption edge detected at 7.42 keV is attributed to K-shell absorption from highly ionized iron in the accretion environment.
@@
To provide a more accurate model description, we further replaced nthcomp with the convolution model thcomp (1996MNRAS.283..193Z; 2019MNRAS.485.2942N; 2020MNRAS.492.5234Z), an updated version of nthcomp that aligns more closely with the actual Monte Carlo spectra from Comptonization than nthcomp does. The model thcomp involves the upscattering fraction (fsc) of photons from any seed emission and gives the low-energy cut-off and the high-energy rollover corresponding to the soft tail of the seed photon and electron temperature of the corona, respectively. As we assumed that the accretion disk supplies the seed photons, we applied the model const*Tbabs*edge*thcomp*diskbb (Model 4). This model also satisfactorily described the persistent spectrum, with = 889/877, and provided similar best-fit parameter values, including absorption edge energies compared to Model 3, suggesting that the source geometry is correctly identified. All the best-fit parameters for the persistent spectral fitting are listed in Table 2. The persistent spectra for the best fit values corresponding to Model 4 is shown in Fig. 3.
@@
We also explored the pre-burst persistent spectra of the source before examining the X-ray burst emissions in detail. For this, we extracted the spectra of the persistent emission from exposures of 600, 600, 1000, 300, 1000, and 1000 s, respectively, just before the onset of bursts B1, B2, B3, B4, B5, and B6, depending on the data availability on the same NuSTAR orbit. The 3 – 60 keV energy spectrum of the pre-burst emission was also successfully fitted with an absorbed thermal Comptonization model along with an absorption edge (Model 3). We note that the best-fit parameter values are similar within error when comparing the values obtained from the persistent emission of the 45 ks exposures described earlier. However, parameters such as electron temperature (k) and edge energy (Eedge) are not well constrained by the pre-burst persistent spectra due to the limited exposures and low statistics (low count rate) before X-ray bursts. We fixed Eedge to 7.42 keV and k to 30 keV which are consistent with our earlier findings obtained from the spectral fit of persistent emission using Model 3 (Table 2). The power-law photon index , seed photon temperature k, absorption optical depth , and the corresponding values obtained from modeling the pre-burst emissions are reported in Table 3.
@@ 4.2 Burst spectral analysis
The burst spectra are shown in Fig. 4. As a best-fit value for all the bursts from Model 6, we obtained the blackbody temperature as k = 1.802 0.027 keV and radius for the blackbody emitting region km. The nthcomp norm is obtained as . This methodology provides us a suitable overall model for the burst spectral analysis, but there may be a significant change of the physical parameters throughout the time-integrated burst spectra. Each burst spectrum comprises of distinct rise, peak and decay phases. There is a rapid change of the physical parameters during each of the phases, particularly during the burst peak. For instance, the temperature changes significantly from the pre-burst to the burst phase, but such rapid variations are averaged out in the time-integrated spectrum. This motivates us to perform the time-resolved analysis for the bursts in Section 4.3 to capture the evolution of the physical parameters. The time-resolved spectral analysis was performed following a similar approach of model adopted by mondal2025nustar for the “clocked burster” GS 1826–24.
@@ 4.3 Time-resolved analysis of the burst spectra
The best-fit parameters of the B1 are shown in Table 4, and the corresponding spectra are shown in lower panel of Fig. 5. The blackbody emitting radius of the neutron star reaches a maximum 3.67 km during S2, where the bbodyrad flux peaks at ergs cm -2 s -1 and then gradually decreases. However, for some segments (S1 and S3) we obtained a low reduced value because of insufficient count rates. The blackbody temperature (k) for B1 reaches a maximum at S2 2.04 keV and then decreases monotonically till S5. The best-fit parameters for time-resolved spectroscopy of B2 are listed in Table 5 and the corresponding spectra are shown in the lower panel of Fig. 6. The bbodyrad flux reaches its peak value during segment S2, after which the bbodyrad flux steadily decreases through S5. The apparent blackbody emitting radius of the neutron star is estimated around 3.09 km during S2, increasing slightly to 3.29 km in S3. A larger radius of 3.71 km is obtained for S5; however, this value is poorly constrained due to limited statistics (0.696), and thus is not considered reliable. The evolution of the blackbody temperature () shows a maximum of keV in S2, followed by a gradual decrease toward S5. The best-fit spectral parameters for B5 across its five time segments are presented in Table 6, with the corresponding spectra and residuals shown in the lower panel of Fig. 7. The blackbody flux reaches its maximum during segment S2 ( erg cm -2 s -1) and subsequently declines in the later segments. The evolution of the blackbody radius closely follows the flux variation, attaining its largest value of approximately 3.5 km during S2, while the blackbody temperature () also attains its highest value at this segment. Across all bursts, the persistent emission shows a pronounced enhancement (), ranging within 1–4 in B1, 1–8 in B2, and 1–7 in B3, suggesting a variable mass accretion rate during the burst evolution (worpel2013evidence).
For the segments B1 S1 and B2 S1, the flux due to persistent emission component (F) exceeds the blackbody flux (Fb) whereas the blackbody flux is expected to dominate until the decay phase of the burst. This may occur due to two possible reasons. Firstly, the time duration of the segments B1 S1 and B2 S1 are larger than the typical rise time of the burst, resulting in the inclusion of the persistent emission effect. If we were able to extract the spectrum for S1 for a time bin size of around 10 s or less than this, we could observe a higher blackbody flux than the persistent one. However, we were unable to perform this because of the limited temporal resolution. Due to the poor counting statistics, for S1 (B1), we have to set the time bin size to 25 s (instead of 20s). Secondly, such behaviour can occur during the sudden onset of a burst, when the ignition may momentarily enhance the accretion rate through Poynting–Robertson drag or due to reprocessing of burst photons in the hot corona. At this early stage, the blackbody component has not yet fully developed, while the persistent Comptonized emission remains comparatively high, leading to a transient phase where the persistent emission exceeds the blackbody flux. Similar short intervals where the persistent flux briefly exceeds the blackbody flux have been reported in previous studies (galloway2008thermonuclear; ji2014hard). The parameter estimates for some other segments, like B1 S5 and B5 S3 yield low reduced values. These low values indicate larger uncertainties in the fits, and the corresponding parameters should therefore be interpreted with caution.
@@ 5 Discussion
We report on the NuSTAR observation of the periodic burster 4U 1323-62, performed on 2024 August 7, for a total exposure of 90 ks. We observed the presence of six thermonuclear X-ray bursts and regular intensity dips in the NuSTAR light curve of this source. The observed burst profiles are very similar to those extracted earlier from the XMM-Newton EPIC PN and AstroSat LAXPC detectors during a persistent interval (boirin2005highly; Bhulla_2020). The observed period ( 2.91-2.98 hr) of the intensity dips is also consistent with the other observations (1985SSRv...40..287V; boirin2005highly; Bhulla_2020). During this observation, the persistent emission is well described by the absorbed thermal Comptonization model in the energy band of 3–60 keV. Our persistent spectral analysis suggests that the Comptonization of soft photons originating from the accretion disc is occurring. The power law index , electron temperature of the corona k, and the seed photon temperatures kTin obtained from the Comptonization model are 1.72 0.01, 18 keV , and 1.58 0.10 keV, respectively. In addition, we detected an absorption edge at an energy of 7.42 keV, indicating the presence of an absorbing medium in the vicinity of the system. Apart from the dip intervals, the persistent flux remains unchanged on our observed timescale. We measured a 3-60 keV persistent flux of F erg cm-2 s-1, which corresponds to a persistent luminosity of Lper 3.321036 erg s-1 for an assumed distance of 10 kpc.
@@
In addition, a 30-70 keV light curve has been plotted to provide another means of diagnosing the burst’s influence on the accretion disk/corona. The light curve for all the bursts, aligned at their start times in both the 30-70 keV and 3-70 keV energy bands, are shown in the lower panel of Fig. 2, where a clear decrease in the count rate during burst interval is evident for all bursts. From the inset plot in the lower panel of Fig. 2, the burst count rate in 30-70 keV hard energy band is observed to vary within 3-4 counts/s at the peak and decrease to less than 3 counts/s during the decay phase. In the 3-70 keV energy band, during the peak emission of the burst, the count rate exceeds 100 counts/s for all bursts except B1 and B4. During the decay phase of the bursts, the count rate decreases to 5-20 counts/s. The drop in the flux in the hard energy band is directly related to the hard X-ray deficits observed during the spectral analysis of type-I X-ray bursts and is considered a key signature of the burst-corona interaction. Thus, the hard X-ray deficit observed in the burst spectra can be explained by the burst-corona interaction (e.g., scattering of the burst photons by the corona). Moreover, during the bursts, the flux from both the thermal and Comptonized components increase, with the thermal photons serving as the seed photons for the Comptonization process. As demonstrated in earlier studies (e.g., ji2014hard), the substantial increase in soft-photon flux leads to enhanced Compton cooling of the corona, thereby reducing the coronal temperature. This observed hard X-ray deficit is a characteristic signature of Type-I thermonuclear bursts.
@@
Except for B4, the burst decay times range from 37.4-50.8 s, classifying them as short bursts ( few minutes). B4 shows distinct properties with , the shortest decay time (29.5 s), and the lowest fluence (0.385 erg cm-2). Such a high typically indicates either accretion-driven emission or incomplete fuel burning (10.1093/mnras/staa2749). In contrast, the subsequent bursts (B5 and B6) with 64–74 are consistent with mixed H/He-fueled events.
@@ 7 Acknowledgements
A note of thanks is extended to the esteemed reviewer for the valuable comments that have improved the quality of this manuscript. We thank NuSTARDAS for ensuring proper extraction of usable data. For the analysis of archival data we deeply acknowledge HEASOFT and CALDB. A token of thanks is extended towards the non-NET fellowship program of Visva-Bharati University for funding the research. We thank IUCAA for its vast repository of knowledge that enlightens curious minds.